how would you write this equation so that y is a function of x?
3+6x=11-4y???

Answer by checkley77(12844)

3+6x=11-4y
6x=11-3-4y
6x=8-4y
x=(8-4y)6
x=2(4-2y)/6
x=(4-2y)/3 ans.